the dean of admissions in a large university has determined that the scores of the freshman class in a mathematics test are apporoximately normally distributed with a mean 72 and a standard deviation of 4 based of this information,what is the probability that the mean score of a sample of 64 students it at least 73


1
Expert's answer
2022-02-10T16:35:33-0500

Let "X=" the mean score of a sample: "X\sim N(\mu, \sigma^2/n)."

Given "\mu=72, \sigma=4, n=64."


"P(X\ge 73)=1-P(X<73)"

"=1-P(Z<\dfrac{73-72}{4/\sqrt{64}})=1-P(Z<2)"

"\approx 0.02275"